\int _3^6\:|12x-10|dx


1
Expert's answer
2022-03-14T17:16:38-0400

Solution


"\int\limits_3^6 {\left| {12x - 10} \right|} \,dx = 2\int\limits_3^6 {\left| {6x - 5} \right|} \,dx"


Then 


"\int {\left| {6x - 5} \right|} \,dx = \frac{1}{6}\int {\left| u \right|} \,du\\
 = \frac{1}{6}\frac{{u\left| u \right|}}{2} + C\\
 = \frac{{u\left| u \right|}}{{12}} + C\\
 = \frac{{\left( {6x - 5} \right)\left| {6x - 5} \right|}}{{12}} + C"


Therefore, 


"\begin{array}{l}
\int\limits_3^6 {\left| {12x - 10} \right|} \,dx = \left[ {2 \times \frac{{\left( {6x - 5} \right)\left| {6x - 5} \right|}}{{12}}} \right]_3^6\\
 = \frac{1}{6}\left[ {\left( {6x - 5} \right)\left| {6x - 5} \right|} \right]_3^6\\
 = \frac{1}{6}\left[ {\left( {6\left( 6 \right) - 5} \right)\left| {6\left( 6 \right) - 5} \right| - \left( {6\left( 3 \right) - 5} \right)\left| {6\left( 3 \right) - 5} \right|} \right]\\
 = \frac{1}{6}\left[ {\left( {31} \right)\left| {31} \right| - \left( {18} \right)\left| {18} \right|} \right]\\
 = 132
\end{array}"